The University of Nebraska football stadium is the third largest city in the state of Nebraska on game days. The stadium has sold out every game since the late 1960s. The seating capacity is about 80,000 fans. Assume the stadium sells out all six home games before the season begins, and the athletic department collects $38.4 million in ticket sales.
   

Required:

  1. 1. What is the average price per season ticket and average price per individual game ticket sold?

  2. 2. & 3. Record the advance collection of $38.4 million in ticket sales and the revenue earned after the first home game is completed.
Answer #1

Answer:

1.

Seating capacity = 80,000

No. Of matches = 6

Total no of audience = 4,80,000

Total collection = $ 38.4 million

Average price per season = $ mill 38.4/80, 000 = 480$

Avg price individual game ticket= 38.4/480000= 80$

Journal entries :

Bank/ cash a/c Dr $38.4million

To advance a/c $ 38.4 million

Advance a/c Dr $ 6.4 million

To sales a/c $6.4 m